### v2

# New Sorting Algorithm Wave Sort (W-Sort)

Abstract.
Modern comparison sorts like quicksort suffer from performance inconsistencies due to suboptimal pivot selection, leading to $O(N^{2})$ worst-case complexity, while in-place merge sort variants face challenges with data movement overhead. We introduce Wave Sort, a novel in-place sorting algorithm that addresses these limitations through a dynamic pivot selection strategy. Wave Sort iteratively expands a sorted region and selects pivots from this growing sorted portion to partition adjacent unsorted data. This approach ensures robust pivot selection irrespective of dataset size, guarantees a logarithmic recursion stack depth, and enables efficient in-place sorting. Our analysis shows a worst-case comparison complexity bounded by $O(N(\log N)^{2})$ with a small constant factor. Experimental results demonstrate that Wave Sort requires significantly fewer comparisons than quicksort on average (approximately 24% less) and performs close to the theoretical minimum, while also incorporating adaptive techniques for efficient handling of presorted sequences. Wave Sort offers a compelling alternative for applications demanding consistent, predictable, and in-place sorting performance.
Sorting algorithms, In-place sorting algorithms, Analysis of algorithms
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Theory of computation Sorting and searching
1. Introduction
The primary cause of inconsistent performance in modern quicksort implementations (Sedgewick and Wayne, 2011a) is lopsided partitions resulting from suboptimal pivot values. This can lead to a significant increase in both the number of comparisons and the depth of the recursion stack, degrading performance from an average of $O(NlogN)$ to a worst-case of $O(N^{2})$. Due to the computational cost associated with selecting a perfectly balanced pivot, current implementations often employ heuristic methods, such as ”median of three”(Sedgewick and Wayne, 2011a). While these methods attempt to choose a pivot from a representative sample, their efficacy can diminish as the dataset size increases because the sample size remains constant and does not scale accordingly. To ensure robust pivot selection irrespective of dataset size, the selection process would ideally need to be dynamic, considering an increasingly larger portion of the data. However, this approach typically increases computational overhead.
Standard merge sort(Sedgewick and Wayne, 2011b) guarantees a comparison time of $O(NlogN)$ and logarithmic stack depth, but it requires additional linear space $O(N)$ and involves $O(NlogN)$ data movements. While in-place merge sort variants(Kim and Kutzner, 2003) eliminate the need for additional space, the complexity of comparison and element movement often increases with higher constant factors.(Huang and Langston, 1988)
We introduce wave sort, a divide-and-conquer, in-place sorting algorithm designed to address the aforementioned shortcomings of quicksort and merge sort. The algorithm commences by sorting an initial small portion of the data. It then utilizes the median value of this sorted portion as the pivot to partition and sort progressively larger segments of the data until the entire dataset is sorted. This strategy effectively expands the algorithm’s ”frame of reference” for selecting a balanced pivot by leveraging already sorted sections of the array. Adaptive sorting is integrated to handle presorted sequence. It is also easy to effectively trade off comparison for element movement. While leapfrogging samplesort(Albacea, 1995) and some variations(Chen, 2007) employ a similar concept, their implementations often lack efficient block swap mechanisms, potentially leading to extra element moves. Furthermore, applying adaptive sorting techniques to these existing variations can be challenging.(Chen, 2006)
Experimental work indicates that, on average, wave sort uses significantly fewer comparisons than quicksort (approximately 24% less) and a comparable number of element moves. The worst-case comparison complexity for wave sort is $O(N(logN)^{2})$, notably with a small constant factor (approximately $\frac{1}{4}$). Adaptive sorting techniques are integrated into wave sort to efficiently handle partially or fully presorted sequences. Additionally, the recursion stack depth is inherently restricted to logarithmic depth. Unlike introsort(Lammich, 2020), wave sort achieves this without requiring explicit monitoring of the stack depth.
2. Core Algorithm
The algorithm is named wave sort because its active working set dynamically expands and contracts during execution, resembling the motion of a wave as it processes the data. Wave sort comprises two primary operations: up-wave and down-wave. The up-wave operation is responsible for expanding the currently sorted region to incorporate an adjacent unsorted portion of the data. Once this expanded working set (containing both the established sorted part and the newly included unsorted part) is defined, it is passed to the down-wave operation for sorting. The down-wave operation then sorts the unsorted portion within this working set using a divide-and-conquer approach. Crucially, pivots for partitioning the unsorted data are selected from the already sorted portion of the working set.
The overall process is iterative:
  1. (1)
    An up-wave phase identifies and incorporates new unsorted data adjacent to the existing sorted region.
  2. (2)
    A subsequent down-wave phase sorts this newly included unsorted data, thereby enlarging the overall sorted region. This cycle repeats until the entire sequence is sorted.
During a down-wave operation:
  • If the unsorted portion within the current working set becomes fully sorted, the down-wave operation for that set is complete. The process may then loop back to the up-wave operation to further expand the now larger sorted region.
  • If the sorted portion (which acts as the pivot provider) is ”exhausted” (all its elements have been used as pivots, or its length is too short for effective partitioning) before the current unsorted portion is fully sorted, the remaining unsorted data is handled by initiating a new up-wave operation. This new operation will establish a new sorted base to continue sorting the remainder.
To simplify the description, and without loss of generality, the following assumptions are made:
  1. (1)
    The array is to be sorted in ascending order (i.e., from the least value to the greatest, with the least value ultimately residing at the leftmost position of the sorted array).
  2. (2)
    The sorting process initiates from the rightmost end. An initial small segment at this end is sorted first, and then the ”wave” of sorting progresses towards the other end.
2.1. Driver function
This function serves as the initial entry point, handling the base case for sequences with fewer than two elements by simply returning. For sequences of two or more elements, it delegates the core sorting logic to the core algorithm.
1type seq []int
2
3func (s seq) WaveSort() {
4 if len(s) < 2 {
5 return
6 }
7 s.upwave(0, len(s)-1)
8}
Listing 1: Driver function
2.2. up-wave
Wave sort commences by establishing an initial sorted segment within the array. Analogous to dynamic programming, this establishes a base case or solves an initial ”subproblem.” This initial segment can be as small as a single element (which is inherently considered sorted) or a larger portion sorted using a conventional algorithm.
Following the establishment of this initial sorted segment, the up-wave operation takes charge. Its primary role is to iteratively expand the bounds of the active working set to incorporate adjacent unsorted data. After each expansion, the up-wave operation passes the newly defined working set (comprising the current sorted region and the newly included unsorted data) to the down-wave operation for sorting.
Once the down-wave completes sorting this set, thereby enlarging the overall sorted region, the up-wave operation re-engages to expand the working set further. This cycle continues until the entire array is sorted. The up-wave thus serves as the main driver and entry point for the overall wave sort process. The mechanics of the up-wave process are illustrated in Fig. 1, and Listing 2.
1func (s seq) upwave(start, end int) {
2 if start == end {
3 return
4 }
5 sortedStart := end
6 sortedLength := 1
7 leftBound := end - 1
8 length := end - start + 1
9 for {
10 s.downwave(leftBound, sortedStart, end)
11 sortedStart = leftBound
12 sortedLength = end - sortedStart + 1
13 if length < sortedLength<<2 {
14 break
15 }
16 leftBound = end - sortedLength<<1 + 1
17 }
18 s.downwave(start, sortedStart, end)
19}
Listing 2: Up-wave Algorithm
Figure 1. up-wave
2.3. down-wave
Before diving into the details of the down-wave, it is necessary to introduce two sub-processes: partitioning and block swapping.
2.3.1. Partitioning
Partitioning splits the sequence into two parts: the left part contains elements less than the pivot, and the right part contains elements greater than the pivot. The function returns the index of the split point, which lies within the greater part.
Initialize pointer i to one position before the first element, and pointer j to one position after the last element. Let p be the pivot. Increment i until the element at index i is greater than or equal to the pivot. If i == j, return i as the middle index m. Then decrement j until the element at index j is less than or equal to the pivot. If j == i, return i as the middle index m. Swap the elements at indices i and j. Repeat this process. The partitioning process is illustrated in Fig. 2 and Listing  3.
1func (s seq) partition(l, r, p int) int {
2 i, j := l-1, r
3 for {
4 for {
5 i++
6 if i == j {
7 return i
8 }
9 if s.less(p, i) {
10 break
11 }
12 }
13 for {
14 j--
15 if j == i {
16 return i
17 }
18 if s.less(j, p) {
19 break
20 }
21 }
22 s.swap(i, j)
23 }
24}
Listing 3: Partitioning Algorithm
Figure 2. Partitioning
2.3.2. Block swapping
Partitioning divides the unsorted portion into left and right blocks, separated by index m. The sorted portion is also split into left and right blocks, separated by the pivot. In the block swapping process, the right block of the unsorted portion (which becomes the left block in the swap) is exchanged with the left block of the sorted portion (which becomes the right block in the swap).
If the left block is shorter than the right block, define its length as $ll=r-m$, where r is the index of the first element in the right block, and m is the index of the first element in the left block. Set $nm=p-ll+1$, and initialize $j=m$. Store the element at index m in a temporary variable. Move the element at index $j+ll$ to index j, then update j to $j+ll$. If $j\geq nm$, replace the element at index $j+ll$ with the one at index $j-nm+m$. Repeat this process for $p-m+1$ iterations, as shown in Listing  4.
1func (s seq) blockSwap_sl(m, p, ll int) {
2 tmp := s[m]
3 init := m
4 j := m
5 nm := p - ll + 1
6 var k int
7 for range p - m + 1 {
8 if j >= nm {
9 k = j - nm + m
10 if k == init {
11 init++
12 s[j] = tmp
13 j = init
14 tmp = s[j]
15 continue
16 }
17 } else {
18 k = j + ll
19 }
20 s[j] = s[k]
21 j = k
22 }
23}
Listing 4: Short left
If the left block is not shorter than the right one, swap elements between the right and left blocks one by one. Only a portion of the unsorted block, equal in size to the sorted block, is moved, as shown in Listing  6.
1func (s seq) blockSwap_sr(m, r, p int) {
2 i := m
3 tmp := s[I]
4 j := r
5 for j < p {
6 s[i] = s[j]
7 i++
8 s[j] = s[i]
9 j++
10 }
11 s[i] = s[j]
12 s[j] = tmp
13}
Listing 5: Short right
All together, the two scenarios are decided by the function shown in Listing  6.
1func (s seq) blockSwap(m, r, p int) {
2 ll := r - m
3 if ll == 0 {
4 return
5 }
6 lr := p - r + 1
7 if lr <= ll {
8 s.blockSwap_sr(m, r, p)
9 bs += lr << 1
10 return
11 }
12 s.blockSwap_sl(m, p, ll)
13}
Listing 6: blockSwap
In both cases, each element is moved only once, with minimal use of a temporary variable. This ensures maximum efficiency. The block swap process is illustrated in Fig. 3.
Figure 3. block swapping
2.3.3. down-wave process
The down-wave selects the median of the sorted portion as the pivot and performs partitioning on the unsorted portion. This partition returns a middle index $m$, which acts as the delimiter between two subsets. Since the pivot corresponds to the median of the sorted portion, the partition ensures that the elements in the left and right subsets of the unsorted section fall on the correct sides of the pivot in the final sorted array.
Specifically, the elements to the left of $m$ in the unsorted portion, along with those to the left of the pivot in the sorted portion, belong to the left of the pivot in the final output. Conversely, the elements to the right of $m$ in the unsorted portion and to the right of the pivot in the sorted portion belong to the right of the pivot. A block swap is performed to exchange the right subset of the unsorted portion with the left subset of the sorted portion. This results in two new sections, each composed of an unsorted part on the left and a sorted part on the right.
The same partition and block swap operations are recursively applied to these new sections, progressively forming smaller subsections with the same structure. This continues until either the sorted or unsorted portion is fully exhausted. If the sorted portion is exhausted, an up-wave can be initiated to process the remaining unsorted portion, starting from its rightmost elements. If the unsorted portion is exhausted, the sorting process is complete. The down-wave procedure is illustrated in Fig. 4 and Listing  7.
1func (s seq) downwave(start, sortedStart, end int) {
2 if sortedStart-start == 0 {
3 return
4 }
5 p := sortedStart + (end-sortedStart)/2
6 m := s.partition(start, sortedStart, p)
7 if m == sortedStart {
8 if p == sortedStart {
9 s.upwave(start, sortedStart-1)
10 return
11 }
12 s.downwave(start, sortedStart, p-1)
13 return
14 }
15 s.blockSwap(m, sortedStart, p)
16 if m == start {
17 if p == sortedStart {
18 s.upwave(m+1, end)
19 return
20 }
21 p++
22 s.downwave(m+p-sortedStart, p, end)
23 return
24 }
25 if p == sortedStart {
26 s.upwave(start, m-1)
27 s.upwave(m+1, end)
28 return
29 }
30 s.downwave(start, m, m+p-sortedStart-1)
31 s.downwave(m+p-sortedStart+1, p+1, end)
32}
Listing 7: Down-wave
Figure 4. down wave
3. Analysis
[…]
$p(i)=\frac{\binom{m}{i-1}\cdot\binom{m}{n-i}}{\binom{n}{2m+1}}$
[…]
$p(i)=\frac{\binom{m-1}{i-1}\cdot\binom{m}{n-i}}{\binom{n}{2m}}$
The probability of evenly dividing the unsorted portion is greater than $\frac{1}{n}$, as is the case in quicksort. Therefore, the average number of comparisons is expected to be closer to the best case than in quicksort. To simplify the analysis, assume the expansion ratio of the up-wave is 2; that is, each up-wave step doubles the size of the working set from the sorted portion. In the best case, the total number of comparisons for the largest up-wave working set is bounded by $C_{0}$:
$C_{0}=\frac{n}{2}+\left(\frac{n}{4}+\frac{n}{4}\right)+\cdots=\frac{n}{2}\log_{2}\frac{n}{2}$
The difference is in each division step of the down-wave, the pivot is removed from the sorted subset. The comparisons for the next smaller working set are bounded by $C_{1}$:
$C_{1}=\frac{n}{4}\log_{2}\frac{n}{4}$
Thus, the total comparison is less than $C_{b}$:
$C_{b}=C_{0}+C_{1}+...=nlog_{2}n-(n-1)$
In the worst case, each comparison provides the minimum information (lowest entropy). Assuming a sorted sequence for analysis, the total number of comparisons is denoted by $C_{w}$:
$C_{w}=\frac{n}{4}\left(\log_{2}^{2}n-\log_{2}\frac{n}{4}\right)$
The maximum stack depth is approximately $\log_{2}n$. In the worst case, the stack depth is less than:
$\frac{1}{2}\left(\log_{2}^{2}n+\log_{2}n\right)$
4. Experimental Results
Experimental code was run on randomly shuffled sequences and compared to quicksort. Results demonstrate that the basic wave sort algorithm performs close to the theoretical minimum comparison limit of $log_{2}(n!)$. Across multiple experiments, wave sort required approximately $1\%$ more comparisons than this theoretical limit on average. This efficiency gap decreases as sequence length increases (See table 1).
sequence size, $n$ wave sort quick sort $\log(n!)$ $C_{b}$
$2^{20}$ 19,680,967 26,070,556 19,458,756 19,922,945
$2^{21}$ 41,459,069 55,091,250 41,014,653 41,943,041
$2^{22}$ 87,113,082 115,662,060 86,223,598 88,080,385
$2^{23}$ 182,613,766 243,388,480 180,835,793 184,549,377
Table 1. Average Comparison
Though wave sort requires fewer comparisons than quicksort, it performs more data swaps and requires additional block swaps. In wave sort, each block swap moves each element once and minimizes access to the temporary variable to only once in most cases. Each swap is equivalent to approximately 3 block swaps. The total data movement in basic wave sort is more than twice that of quicksort (See table 2).
sequence size, $n$ wave sort, swap wave sort, block swap quick sort, swap
$2^{20}$ 5,001,848 16,984,185 4,757,387
$2^{21}$ 10,528,024 36,064,105 9,994,835
$2^{22}$ 22,104,628 76,317,658 20,976,731
$2^{23}$ 46,307,301 161,022,256 43,869,539
Table 2. Average Swap
In practice, many sequences encountered are presorted or partially presorted. Wave sort exhibits its worst-case performance when processing fully presorted sequences (See table 3).
sequence size, $n$ wave sort $C_{w}$
$2^{20}$ 100,139,008 100,139,008
$2^{21}$ 221,249,536 221,249,536
$2^{22}$ 486,539,264 486,539,264
$2^{23}$ 1,065,353,216 1,065,353,216
Table 3. Worst Comparison
5. Adaptive Improvement
For sorted or partially sorted sequences, regardless of whether they are reversed, an additional check is incorporated into the up-wave to identify the longest presorted sequence as the sorted portion. For a fully sorted sequence, this results in only $n-1$ comparisons. If the sequence is completely reversed, the number of swaps is reduced to $\frac{n}{2}$. This modification effectively short-circuits the processing of presorted sequences (see Listing  8 and table 4).
1func (s seq) upwave(start, end int) {
2 if start == end {
3 return
4 }
5 sortedStart := s.preSorted(start, end)
6 sortedLength := end - sortedStart + 1
7 leftBound := end - sortedLength<<1 + 1
8 length := end - start + 1
9 for leftBound > start {
10 s.downwave(leftBound, sortedStart, end)
11 sortedStart = leftBound
12 sortedLength = end - sortedStart + 1
13 if length < sortedLength<<2 {
14 break
15 }
16 leftBound = end - sortedLength<<1 + 1
17 }
18 s.downwave(start, sortedStart, end)
19}
20
21func (s seq) preSorted(start, end int) int {
22 du := false
23 i := end
24 for i > start {
25 if du {
26 if s.less(i, i-1) {
27 return I
28 }
29 i--
30 continue
31 }
32 if s.less(i-1, i) {
33 if i == end {
34 du = true
35 i--
36 continue
37 }
38 s.reverse(i, end)
39 return i
40 }
41 i--
42 }
43 if du {
44 return start
45 }
46 s.reverse(start, end)
47 return start
48}
49
50func (s seq) reverse(start, end int) {
51 i := start
52 j := end
53 for range (end - start + 1) / 2 {
54 s.swap(i, j)
55 i++
56 j--
57 }
58}
Listing 8: Adaptive Wave Sort
sequence size, $n$ wave sort adaptive reverse swap
$2^{20}$ 1,048,575 524,288
$2^{21}$ 2,097,151 1,048,576
$2^{22}$ 4,194,303 2,097,152
$2^{23}$ 8,388,607 4,194,304
Table 4. Adaptive Presorted
As the adaptive approach adds one more step to check the boundary of the sorted sequence, the number of comparisons increases slightly. This approach trades off swaps for block swaps. The comparisons increase by approximately $0.6\%$. The number of swaps decreases by approximately $1\%$, while block swaps decrease by approximately $7\%$ (See table 5).
sequence size, $n$ comparison swap block swap
$2^{20}$ 19,824,134 4,972,230 15,569,767
$2^{21}$ 41,729,429 10,458,356 33,523,711
$2^{22}$ 87,669,368 21,975,151 70,442,223
$2^{23}$ 183,615,126 45,971,847 152,140,092
Table 5. Adaptive Average
6. Trade off Comparison for Swap
The average comparison count of wave sort is quite close to the theoretical limit, leaving only a small margin for improvement. In practice, comparisons usually cost more than swaps on complex data structures, as comparisons require examining the content of the data structure while swaps only require exchanging pointers to the content. Therefore, fewer comparisons are generally preferred. However, in some cases, trading comparisons for swaps may be advantageous.
Since wave sort performs more than twice the swaps compared to quicksort, there is room for optimization if such trade-offs prove beneficial. The basic strategy is to increase the expansion rate in the up-wave so that partitioning works more effectively to reduce swaps. In both down-wave and up-wave processes, sorting short sequences with insertion sort can replace swaps with block swaps for greater efficiency. The smaller sorted portions also reduce the number of block swaps required. An implementation showing these trade-offs is present in Listing  9
1func (s seq) upwave(start, end int) {
2 length := end - start + 1
3 if length == 1 {
4 return
5 }
6 if length <= 8 {
7 s.insertSort(start, end)
8 return
9 }
10 sortedStart := s.preSorted(start, end)
11 sortedLength := end - sortedStart + 1
12 sortedLength <<= 4
13 leftBound := end - sortedLength + 1
14 for leftBound > start {
15 s.downwave(leftBound, sortedStart, end)
16 sortedStart = leftBound
17 sortedLength = end - sortedStart + 1
18 if length <= sortedLength<<6 {
19 break
20 }
21 sortedLength <<= 4
22 leftBound = end - sortedLength + 1
23 }
24 s.downwave(start, sortedStart, end)
25}
26
27func (s seq) downwave(start, sortedStart, end int) {
28 unsortedLength := sortedStart - start
29 if unsortedLength == 0 {
30 return
31 }
32 lr := end - sortedStart
33 if lr == 0 {
34 if unsortedLength < 8 {
35 s.insertSort(start, end)
36 return
37 }
38 s.upwave(start, end)
39 return
40 }
41 p := sortedStart + lr/2
42 ...
43}
44
45func (s seq) insertSort(l, r int) {
46 sortedIndex := l + 1
47 for sortedIndex <= r {
48 i := sortedIndex
49 for i > l {
50 if s.less(sortedIndex, i-1) {
51 i--
52 continue
53 }
54 break
55 }
56 if i < sortedIndex {
57 tmp := s[sortedIndex]
58 for j := sortedIndex; j > i; j-- {
59 s[j] = s[j-1]
60 }
61 s[i] = tmp
62 }
63 sortedIndex++
64 }
65}
Listing 9: Trade-off Comparison
A typical trade-off result is shown in Table 6 with the implementation demonstrated in the code above. One comparison trades for approximately 4 swaps plus 80 block swaps, which is equivalent to approximately 31 swaps (See table 6). More detailed mathematical analysis is required to determine the optimal expansion rate. Generally, the longer the sorted portion, the larger the expansion rate should be, based on the probability equation provided above.
sequence size, $n$ comparison swap block swap
$2^{20}$ 19,853,004 4,467,539 4,459,874
$2^{21}$ 41,996,088 9,462,760 7,691,231
$2^{22}$ 88,300,820 19,888,658 14,535,769
$2^{23}$ 185,437,750 41,802,470 26,231,740
Table 6. Trade off Comparison for Swap
7. Conclusion
In this paper, we introduced Wave Sort, a novel divide-and-conquer in-place sorting algorithm designed to address key limitations found in widely used comparison sorts. By employing a dynamic pivot selection strategy that leverages an expanding sorted region within the array, Wave Sort overcomes the pivot inconsistency issues that can lead to $O(N^{2})$ worst-case performance in quicksort. Unlike standard merge sort, Wave Sort achieves sorting in-place, avoiding the need for $O(N)$ auxiliary space, and ensures a consistent, low logarithmic recursion stack depth without the need for explicit monitoring, presenting an advantage over algorithms like Introsort in memory-constrained environments.
Our analysis indicates a worst-case comparison complexity of $O(N(\log N)^{2})$ with a small constant factor ($\approx\frac{1}{4}$). Experimental evaluation on randomly shuffled data demonstrated that Wave Sort achieves an average comparison count significantly lower than quicksort (approximately 24% less) and remarkably close to the theoretical minimum of $O(N\log N)$, aligning with our analytical expectations. Furthermore, the algorithm is readily adaptable to handle partially or fully presorted and reversed sequences with high efficiency, achieving linear time complexity in these best-case scenarios through the integrated adaptive improvement. While the basic implementation involves more data movement operations than quicksort, we showed that the algorithm’s structure allows for effective trade-offs to reduce swaps by increasing comparison operations, depending on the relative costs in a given application context.
Wave Sort presents a compelling alternative for scenarios requiring efficient in-place sorting with predictable performance characteristics. Its combination of guaranteed logarithmic stack depth, favorable average comparison performance, and bounded worst-case behavior positions it as a strong candidate for replacing algorithms like in-place merge sort variants and Introsort in standard library implementations, especially where robust and consistent sorting is paramount.
Future work includes conducting a more detailed formal analysis to determine the optimal parameters for trading comparisons against element moves across various data types and hardware architectures, thereby maximizing practical performance. Further research could also explore variations in the initial sorted segment size, alternative median or pivot selection strategies from the sorted portion, potential parallelization of the algorithm, and its applicability to related sorting problems or data structures. Wave Sort represents a promising direction for advancing the state of the art in in-place comparison sorting by effectively balancing theoretical elegance with practical performance considerations.
Acknowledgements.
For my father, Wei Wei, who motivated me to write this, and Professor Ghosal for his very helpful stats class.